The sum of the interior angles of an n-sided regular polygon is 720 degree. Find the size of each exterior angle of the polygon.
1
Expert's answer
2020-04-17T17:23:30-0400

for a regular polygon, interior angle is given by,


"interior\ angle=180^\circ-\frac{360^\circ}{n}\\"

Since there are n edges,

The sum of the interior angles ="interior\ angle *n=(180^\circ-\frac{360^\circ}{n})n\\"

In this question sum of interior angles is "720^\circ"

Therefore number of sided(n),

"720^\circ=(180^\circ-\frac{360^\circ}{n})n\\
4=(1-\frac{2}{n})n\\
\frac{4}{n}=1-\frac{2}{n}\\
\frac{6}{n}=1\\
n=6"

Therefore size of interior angle ="\frac{720^\circ}{6}=120^\circ"

"\therefore Each\ exterior\ angle=180^\circ-120^\circ\\
\therefore Each\ exterior\ angle=\bold{60^\circ }"


Also in another way,

"\therefore Each\ exterior\ angle=\frac{360^\circ}{6}=\bold{60^\circ}"